what is the value of x?
4^20+4^20=2^x

Answer by stanbon(57962)

what is the value of x?
4^20+4^20=2^x
-----------
2(4^20) = 2^x
2(2^2*20) = 2^x
2*2^40 = 2^x
2^41 = 2^x
x = 41
===========
Cheers,
Stan H.